# CRAFT: Calibrated Reasoning with Answer-Faithful Traces via Reinforcement Learning for Multi-Hop Question Answering

### v1

@@ C.1 Training Dynamics
| EM | F1 | Faithfulness | EM | F1 | Faithfulness | EM | F1 | Faithfulness | ||
| Open-source | ||||||||||
| Qwen2.5-0.5B Team [2024] | Pre-Trained | 0.25 | 1.72 | 0.85 | 1.80 | 7.31 | 6.40 | 1.85 | 8.81 | 4.60 |
| Qwen2.5-1.5B Team [2024] | Pre-Trained | 2.15 | 6.98 | 15.55 | 7.30 | 16.58 | 35.20 | 10.05 | 17.10 | 21.35 |
| Qwen2.5-3B Team [2024] | Pre-Trained | 17.70 | 26.19 | 38.85 | 38.85 | 51.77 | 69.70 | 34.30 | 43.93 | 62.00 |
| Qwen2.5-7B Team [2024] | Pre-Trained | 38.95 | 49.51 | 68.70 | 57.75 | 72.10 | 90.65 | 60.65 | 69.82 | 84.25 |
| Qwen3-4B Team [2025] | Pre-Trained | 29.75 | 39.04 | 57.40 | 54.60 | 70.04 | 88.60 | 56.20 | 66.71 | 86.80 |
| Qwen3-30B Team [2025] | Pre-Trained | 43.80 | 53.59 | 72.65 | 61.05 | 75.94 | 93.80 | 63.95 | 73.70 | 90.30 |
| API | ||||||||||
| GPT-5-mini OpenAI [2025] | API | 52.45 | 67.83 | 89.45 | 60.98 | 79.77 | 98.20 | 69.60 | 80.28 | 97.25 |
| Deepseek-V3.2 DeepSeek-AI [2025] | API | 52.75 | 67.04 | 83.40 | 62.90 | 80.00 | 97.10 | 68.90 | 78.06 | 97.65 |
| Gemini-2.5-Flash Google DeepMind [2025] | API | 51.95 | 63.53 | 82.65 | 64.00 | 79.59 | 96.00 | 70.15 | 79.45 | 96.85 |
| SFT | ||||||||||
| Qwen2.5-0.5B Team [2024] | Trained | 1.10 | 5.92 | 5.35 | 12.80 | 21.39 | 17.45 | 22.45 | 27.18 | 17.95 |
| Qwen2.5-1.5B Team [2024] | Trained | 9.65 | 17.17 | 15.55 | 36.15 | 47.65 | 45.10 | 36.00 | 43.19 | 38.00 |
| Qwen2.5-3B Team [2024] | Trained | 22.15 | 31.93 | 43.85 | 51.40 | 65.75 | 79.90 | 53.40 | 61.93 | 71.90 |
| Qwen2.5-7B Team [2024] | Trained | 35.60 | 47.66 | 61.35 | 59.25 | 73.68 | 88.25 | 65.40 | 74.33 | 85.80 |
| Ours (RL) | ||||||||||
@@
| EM | F1 | Faithfulness | EM | F1 | Faithfulness | EM | F1 | Faithfulness | ||
| Open-source | ||||||||||
| Qwen2.5-0.5B Team [2024] | Pre-Trained | 0.35 | 2.69 | 4.85 | 3.00 | 7.51 | 14.10 | 5.05 | 9.98 | 8.00 |
| Qwen2.5-1.5B Team [2024] | Pre-Trained | 4.85 | 9.78 | 16.65 | 26.80 | 36.98 | 44.65 | 21.65 | 28.33 | 31.15 |
| Qwen2.5-3B Team [2024] | Pre-Trained | 21.95 | 32.30 | 41.15 | 48.30 | 61.44 | 76.05 | 41.95 | 50.67 | 60.80 |
| Qwen2.5-7B Team [2024] | Pre-Trained | 38.35 | 49.29 | 76.60 | 59.30 | 74.40 | 93.75 | 60.80 | 70.20 | 86.15 |
| Qwen3-4B Team [2025] | Pre-Trained | 38.50 | 47.61 | 77.55 | 60.70 | 76.06 | 95.60 | 65.40 | 74.61 | 93.90 |
| Qwen3-30B Team [2025] | Pre-Trained | 46.75 | 56.34 | 84.95 | 63.35 | 78.13 | 97.10 | 70.35 | 79.04 | 95.95 |
| API | ||||||||||
| GPT-5-mini OpenAI [2025] | API | 52.05 | 68.10 | 93.15 | 60.85 | 79.83 | 98.05 | 70.05 | 81.09 | 97.10 |
| Deepseek-V3.2 DeepSeek-AI [2025] | API | 53.20 | 67.96 | 88.25 | 61.55 | 78.67 | 98.25 | 70.35 | 79.35 | 98.55 |
| Gemini-2.5-Flash Google DeepMind [2025] | API | 51.25 | 62.92 | 90.20 | 63.75 | 79.65 | 97.20 | 70.60 | 79.41 | 97.80 |
| SFT | ||||||||||
| Qwen2.5-0.5B Team [2024] | Trained | 0.50 | 4.62 | 13.50 | 11.75 | 21.27 | 37.40 | 23.15 | 28.68 | 26.85 |
| Qwen2.5-1.5B Team [2024] | Trained | 7.25 | 12.85 | 16.40 | 33.25 | 44.48 | 51.95 | 36.30 | 43.53 | 42.70 |
| Qwen2.5-3B Team [2024] | Trained | 23.60 | 33.10 | 50.15 | 50.25 | 64.37 | 83.85 | 55.90 | 64.77 | 74.10 |
| Qwen2.5-7B Team [2024] | Trained | 31.80 | 42.71 | 66.60 | 57.90 | 72.43 | 89.35 | 63.15 | 72.00 | 82.25 |
| Ours (RL) | ||||||||||
@@
| Open-source | |||||||
| Qwen2.5-0.5B Team [2024] | Pre-Trained | 0.30 | 3.20 | 5.45 | 12.18 | 6.40 | 12.37 |
| Qwen2.5-1.5B Team [2024] | Pre-Trained | 10.20 | 17.58 | 35.95 | 47.77 | 26.05 | 33.29 |
| Qwen2.5-3B Team [2024] | Pre-Trained | 23.20 | 33.31 | 49.30 | 62.58 | 41.45 | 50.09 |
| Qwen2.5-7B Team [2024] | Pre-Trained | 31.55 | 41.97 | 53.40 | 67.92 | 50.60 | 59.29 |
| Qwen3-4B Team [2025] | Pre-Trained | 39.55 | 49.23 | 60.40 | 75.32 | 57.70 | 67.43 |
| Qwen3-30B Team [2025] | Pre-Trained | 49.30 | 59.33 | 63.05 | 78.33 | 64.70 | 73.71 |
| API | |||||||
| GPT-5-mini OpenAI [2025] | API | 51.10 | 64.43 | 63.63 | 80.19 | 65.28 | 76.53 |
| Deepseek-V3.2 DeepSeek-AI [2025] | API | 58.10 | 70.97 | 64.15 | 79.91 | 67.45 | 76.66 |
| Gemini-2.5-Flash Google DeepMind [2025] | API | 54.80 | 66.64 | 65.68 | 80.79 | 71.85 | 80.59 |
| SFT | |||||||
| Qwen2.5-0.5B Team [2024] | Trained | 0.15 | 0.85 | 0.80 | 2.85 | 2.80 | 4.70 |
| Qwen2.5-1.5B Team [2024] | Trained | 10.05 | 17.23 | 34.10 | 45.92 | 34.30 | 40.92 |
| Qwen2.5-3B Team [2024] | Trained | 19.60 | 28.60 | 46.60 | 60.23 | 50.40 | 57.75 |
| Qwen2.5-7B Team [2024] | Trained | 27.00 | 37.82 | 53.30 | 68.12 | 56.15 | 63.79 |
| Ours (RL) | |||||||
@@
| HotpotQA | 0.0 | 99.6(+99.6) | 0.1 | 99.8(+99.7) | 0.1 | 97.5(+97.4) | 0.8 | 99.8(+99.0) | 5.7 | 0.1(-5.6) | |
| 2WikiMQA | 0.0 | 99.5(+99.5) | 3.2 | 99.7(+96.5) | 0.0 | 92.5(+92.5) | 2.3 | 99.8(+97.5) | 16.5 | 0.1(-16.4) | |
| MuSiQue | 0.0 | 95.8(+95.8) | 0.0 | 98.2(+98.2) | 0.0 | 92.8(+92.8) | 0.0 | 98.5(+98.5) | 4.9 | 0.0(-4.9) | |
| Qwen2.5-1.5B | |||||||||||
@@
| HotpotQA | 0.0 | 99.5(+99.5) | 2.9 | 98.9(+96.0) | 0.1 | 99.2(+99.1) | 1.0 | 92.9(+91.9) | 0.7 | 100.0(+99.3) | |
| 2WikiMQA | 0.0 | 99.4(+99.4) | 0.8 | 99.2(+98.4) | 0.0 | 99.5(+99.5) | 1.4 | 91.2(+89.8) | 1.4 | 100.0(+98.6) | |
| MuSiQue | 0.1 | 99.1(+99.0) | 2.5 | 87.4(+84.9) | 0.6 | 99.7(+99.1) | 1.2 | 61.6(+60.4) | 0.2 | 99.9(+99.7) | |
| Qwen2.5-3B | |||||||||||
@@
| HotpotQA | 14.8 | 98.2(+83.4) | 55.2 | 99.2(+44.0) | 4.3 | 99.4(+95.1) | 27.6 | 99.2(+71.6) | 32.8 | 100.0(+67.2) | |
| 2WikiMQA | 7.7 | 99.0(+91.3) | 33.8 | 99.4(+65.6) | 6.3 | 99.1(+92.8) | 8.2 | 98.8(+90.6) | 23.9 | 100.0(+76.1) | |
| MuSiQue | 17.2 | 93.5(+76.3) | 61.8 | 97.6(+35.8) | 8.7 | 95.6(+86.9) | 10.8 | 96.3(+85.5) | 68.1 | 100.0(+31.9) | |
| Qwen2.5-7B | |||||||||||
@@
| HotpotQA | 96.1 | 100.0(+3.9) | 99.5 | 100.0(+0.5) | 95.7 | 99.1(+3.4) | 49.5 | 100.0(+50.5) | 11.2 | 100.0(+88.8) | |
| 2WikiMQA | 95.4 | 99.1(+3.7) | 99.3 | 100.0(+0.7) | 96.0 | 97.0(+1.0) | 72.9 | 100.0(+27.1) | 9.6 | 100.0(+90.4) | |
| MuSiQue | 90.9 | 98.9(+8.0) | 98.9 | 100.0(+1.1) | 91.6 | 95.5(+3.9) | 62.5 | 99.4(+36.9) | 25.4 | 100.0(+74.6) | |
C.2 Evaluation Results
@@
CRAFT retains the same four-field structure as CRAFT but with simplified instructions. CRAFT achieves strong performance: 58.35% EM on MuSiQue (+24.0% vs Base), 67.28% on HotpotQA (+12.5% vs Base), and 79.70% on 2WikiMHQA (+23.7% vs Base). Notably, CRAFT surpasses all API models on MuSiQue and 2WikiMHQA while achieving competitive Faithfulness scores (78.35%–95.20%).
@@
CRAFT further simplifies the template while maintaining structured reasoning. CRAFT achieves the highest Faithfulness scores among all templates: 86.70% on MuSiQue, 96.00% on HotpotQA, and 94.10% on 2WikiMHQA. EM scores (56.91%, 66.51%, 79.39%) remain competitive, demonstrating that simplified templates can achieve strong performance when combined with appropriate reward design.
@@
CRAFT represents the minimal template with direct answer generation, lacking structured reasoning traces ( and not applicable). CRAFT achieves 56.25% EM on MuSiQue, 63.25% on HotpotQA, and 67.15% on 2WikiMHQA. A critical observation: CRAFT completely fails on this template (near-zero performance), indicating that smaller models require structured reasoning traces for stable training. This validates the importance of trace structure for model capacity constraints.
CRAFT Template Sensitivity Analysis.
Our experiments reveal a striking template sensitivity pattern unique to the smallest model, with different failure modes for SFT and GRPO training. SFT training collapses on templates CRAFT, CRAFT, and CRAFT (EM 1.5% across benchmarks), achieving meaningful performance only on CRAFT (1.10/12.80/22.45%) and CRAFT (0.50/11.75/23.15%). In contrast, GRPO rescues templates CRAFT–CRAFT but still fails on CRAFT. Specifically, GRPO achieves 7.25/24.80/22.85% on CRAFT, 6.50/20.60/22.65% on CRAFT, 4.75/20.20/20.50% on CRAFT, and 8.45/25.95/24.70% on CRAFT, while CRAFT remains completely broken (0.00/0.00/0.05%).
This dichotomy reveals two key insights: (1) SFT cold-start is fragile for 0.5B models—templates CRAFT/CRAFT with explicit <gold_docs> fields overwhelm supervised imitation, causing training collapse. (2) GRPO’s exploration mechanism overcomes SFT’s brittleness—reinforcement learning enables the model to discover valid reasoning paths even on complex structures (CRAFT/CRAFT), as long as some structural scaffolding exists. (3) Template CRAFT’s complete failure across both training paradigms confirms that 0.5B models fundamentally require explicit reasoning structure—direct answer generation lacks sufficient guidance for stable learning.
The optimal template for 0.5B deployment is CRAFT (<plan>+<reason>+<answer> without <gold_docs>), which provides the best balance: enough structure to guide reasoning without exceeding representational capacity, and compatibility with both SFT initialization and GRPO refinement.
C.3 Output Traces Format Analysis

### v2

@@ Appendix A Implementation Details
| EM | F1 | Faith. | EM | F1 | Faith. | EM | F1 | Faith. | ||
[…]
| Qwen2.5-0.5B | Pre-Trained | 0.25 | 1.72 | 0.85 | 1.80 | 7.31 | 6.40 | 1.85 | 8.81 | 4.60 |
| Qwen2.5-1.5B | Pre-Trained | 2.15 | 6.98 | 15.55 | 7.30 | 16.58 | 35.20 | 10.05 | 17.10 | 21.35 |
| Qwen2.5-3B | Pre-Trained | 17.70 | 26.19 | 38.85 | 38.85 | 51.77 | 69.70 | 34.30 | 43.93 | 62.00 |
| Qwen2.5-7B | Pre-Trained | 38.95 | 49.51 | 68.70 | 57.75 | 72.10 | 90.65 | 60.65 | 69.82 | 84.25 |
| Qwen3-4B | Pre-Trained | 29.75 | 39.04 | 57.40 | 54.60 | 70.04 | 88.60 | 56.20 | 66.71 | 86.80 |
| Qwen3-30B | Pre-Trained | 43.80 | 53.59 | 72.65 | 61.05 | 75.94 | 93.80 | 63.95 | 73.70 | 90.30 |
[…]
| GPT-5-mini | API | 52.45 | 67.83 | 89.45 | 60.98 | 79.77 | 98.20 | 69.60 | 80.28 | 97.25 |
| Deepseek-V3.2 | API | 52.75 | 67.04 | 83.40 | 62.90 | 80.00 | 97.10 | 68.90 | 78.06 | 97.65 |
| Gemini-2.5-Flash | API | 51.95 | 63.53 | 82.65 | 64.00 | 79.59 | 96.00 | 70.15 | 79.45 | 96.85 |
[…]
| Qwen2.5-0.5B | Trained | 1.10 | 5.92 | 5.35 | 12.80 | 21.39 | 17.45 | 22.45 | 27.18 | 17.95 |
| Qwen2.5-1.5B | Trained | 9.65 | 17.17 | 15.55 | 36.15 | 47.65 | 45.10 | 36.00 | 43.19 | 38.00 |
| Qwen2.5-3B | Trained | 22.15 | 31.93 | 43.85 | 51.40 | 65.75 | 79.90 | 53.40 | 61.93 | 71.90 |
| Qwen2.5-7B | Trained | 35.60 | 47.66 | 61.35 | 59.25 | 73.68 | 88.25 | 65.40 | 74.33 | 85.80 |
[…]
@@
| EM | F1 | Faith. | EM | F1 | Faith. | EM | F1 | Faith. | ||
[…]
| Qwen2.5-0.5B | Pre-Trained | 0.35 | 2.69 | 4.85 | 3.00 | 7.51 | 14.10 | 5.05 | 9.98 | 8.00 |
| Qwen2.5-1.5B | Pre-Trained | 4.85 | 9.78 | 16.65 | 26.80 | 36.98 | 44.65 | 21.65 | 28.33 | 31.15 |
| Qwen2.5-3B | Pre-Trained | 21.95 | 32.30 | 41.15 | 48.30 | 61.44 | 76.05 | 41.95 | 50.67 | 60.80 |
| Qwen2.5-7B | Pre-Trained | 38.35 | 49.29 | 76.60 | 59.30 | 74.40 | 93.75 | 60.80 | 70.20 | 86.15 |
| Qwen3-4B | Pre-Trained | 38.50 | 47.61 | 77.55 | 60.70 | 76.06 | 95.60 | 65.40 | 74.61 | 93.90 |
| Qwen3-30B | Pre-Trained | 46.75 | 56.34 | 84.95 | 63.35 | 78.13 | 97.10 | 70.35 | 79.04 | 95.95 |
[…]
| GPT-5-mini | API | 52.05 | 68.10 | 93.15 | 60.85 | 79.83 | 98.05 | 70.05 | 81.09 | 97.10 |
| Deepseek-V3.2 | API | 53.20 | 67.96 | 88.25 | 61.55 | 78.67 | 98.25 | 70.35 | 79.35 | 98.55 |
| Gemini-2.5-Flash | API | 51.25 | 62.92 | 90.20 | 63.75 | 79.65 | 97.20 | 70.60 | 79.41 | 97.80 |
[…]
| Qwen2.5-0.5B | Trained | 0.50 | 4.62 | 13.50 | 11.75 | 21.27 | 37.40 | 23.15 | 28.68 | 26.85 |
| Qwen2.5-1.5B | Trained | 7.25 | 12.85 | 16.40 | 33.25 | 44.48 | 51.95 | 36.30 | 43.53 | 42.70 |
| Qwen2.5-3B | Trained | 23.60 | 33.10 | 50.15 | 50.25 | 64.37 | 83.85 | 55.90 | 64.77 | 74.10 |
| Qwen2.5-7B | Trained | 31.80 | 42.71 | 66.60 | 57.90 | 72.43 | 89.35 | 63.15 | 72.00 | 82.25 |
[…]
@@
[…]
| Qwen2.5-0.5B | Pre-Trained | 0.30 | 3.20 | 5.45 | 12.18 | 6.40 | 12.37 |
| Qwen2.5-1.5B | Pre-Trained | 10.20 | 17.58 | 35.95 | 47.77 | 26.05 | 33.29 |
| Qwen2.5-3B | Pre-Trained | 23.20 | 33.31 | 49.30 | 62.58 | 41.45 | 50.09 |
| Qwen2.5-7B | Pre-Trained | 31.55 | 41.97 | 53.40 | 67.92 | 50.60 | 59.29 |
| Qwen3-4B | Pre-Trained | 39.55 | 49.23 | 60.40 | 75.32 | 57.70 | 67.43 |
| Qwen3-30B | Pre-Trained | 49.30 | 59.33 | 63.05 | 78.33 | 64.70 | 73.71 |
[…]
| GPT-5-mini | API | 51.10 | 64.43 | 63.63 | 80.19 | 65.28 | 76.53 |
| Deepseek-V3.2 | API | 58.10 | 70.97 | 64.15 | 79.91 | 67.45 | 76.66 |
| Gemini-2.5-Flash | API | 54.80 | 66.64 | 65.68 | 80.79 | 71.85 | 80.59 |
[…]
| Qwen2.5-0.5B | Trained | 0.15 | 0.85 | 0.80 | 2.85 | 2.80 | 4.70 |
| Qwen2.5-1.5B | Trained | 10.05 | 17.23 | 34.10 | 45.92 | 34.30 | 40.92 |
| Qwen2.5-3B | Trained | 19.60 | 28.60 | 46.60 | 60.23 | 50.40 | 57.75 |
| Qwen2.5-7B | Trained | 27.00 | 37.82 | 53.30 | 68.12 | 56.15 | 63.79 |
[…]
@@
| HotpotQA | 0.0 | 99.6 | 0.1 | 99.8 | 0.1 | 97.5 | 0.8 | 99.8 | 5.7 | 0.1 | |
| 2WikiMQA | 0.0 | 99.5 | 3.2 | 99.7 | 0.0 | 92.5 | 2.3 | 99.8 | 16.5 | 0.1 | |
| MuSiQue | 0.0 | 95.8 | 0.0 | 98.2 | 0.0 | 92.8 | 0.0 | 98.5 | 4.9 | 0.0 | |
[…]
@@
| HotpotQA | 0.0 | 99.5 | 2.9 | 98.9 | 0.1 | 99.2 | 1.0 | 92.9 | 0.7 | 100.0 | |
| 2WikiMQA | 0.0 | 99.4 | 0.8 | 99.2 | 0.0 | 99.5 | 1.4 | 91.2 | 1.4 | 100.0 | |
| MuSiQue | 0.1 | 99.1 | 2.5 | 87.4 | 0.6 | 99.7 | 1.2 | 61.6 | 0.2 | 99.9 | |
[…]
@@
| HotpotQA | 14.8 | 98.2 | 55.2 | 99.2 | 4.3 | 99.4 | 27.6 | 99.2 | 32.8 | 100.0 | |
| 2WikiMQA | 7.7 | 99.0 | 33.8 | 99.4 | 6.3 | 99.1 | 8.2 | 98.8 | 23.9 | 100.0 | |
| MuSiQue | 17.2 | 93.5 | 61.8 | 97.6 | 8.7 | 95.6 | 10.8 | 96.3 | 68.1 | 100.0 | |
[…]
@@
| HotpotQA | 96.1 | 100.0 | 99.5 | 100.0 | 95.7 | 99.1 | 49.5 | 100.0 | 11.2 | 100.0 | |
| 2WikiMQA | 95.4 | 99.1 | 99.3 | 100.0 | 96.0 | 97.0 | 72.9 | 100.0 | 9.6 | 100.0 | |
| MuSiQue | 90.9 | 98.9 | 98.9 | 100.0 | 91.6 | 95.5 | 62.5 | 99.4 | 25.4 | 100.0 | |
Appendix B Additional Results
This section provides comprehensive supplementary results that expand on the main experiments in §4. We present training dynamics for all trace variants (§B.1), verification of faithfulness reward behavior (§B.2), full evaluation tables for CRAFT–CRAFT (§B.3), format compliance analysis (§B.4), format error patterns (§B.5), and model scale sensitivity (§B.6).
B.1 Training Dynamics for Additional Variants
Figures 5–8 show the training dynamics for CRAFT–CRAFT. Consistent with CRAFT (Figure 3), all templates exhibit smooth convergence with larger models achieving higher total rewards.
Template-specific patterns: (1) CRAFT shows similar dynamics to CRAFT, with all four reward components contributing. (2) CRAFT/CRAFT lack but maintain stable training. (3) CRAFT uses only and ; the 0.5B model shows minimal improvement, foreshadowing its evaluation failure.
B.2 Faithfulness Reward Verification
As noted in §4.4, the faithfulness reward remains low during training, particularly for smaller models. To verify that this reflects optimization dynamics rather than learning failure, we measured the four faithfulness metrics on test samples of the Qwen2.5-1.5B CRAFT model under greedy decoding (temp=0.0). Among test samples with correct answers (EM=1), the four faithfulness metrics achieve 76.85%, 68.95%, 70.15%, and 61.35% respectively, with an arithmetic mean of 69.33%. This indicates that while 69.3% of correct predictions exhibit fully faithful reasoning, 30.7% still fail at least one faithfulness check, demonstrating that answer correctness alone does not guarantee reasoning faithfulness. This test-time value is approximately six times higher than the training-time average of 0.11, confirming that GRPO’s on-policy sampling at high temperature (temp=1.0) generates diverse outputs including many low-quality traces during exploration, while the learned policy concentrates probability mass on high-quality outputs under greedy decoding at test time.
B.3 Full Results for Additional Variants
Tables A–A present the complete evaluation results for CRAFT–CRAFT across all model scales and baselines, complementing the CRAFT results in the main paper (Table 3.3.2).
@@ B.3 Full Results for Additional Variants
CRAFT retains the same structure as CRAFT but removes <plan>. CRAFT achieves strong performance: 58.35% EM on MuSiQue (+24.0% vs Base), 67.28% on HotpotQA (+12.5% vs Base), and 79.70% on 2WikiMHQA (+23.7% vs Base). Notably, CRAFT surpasses all API models on MuSiQue and 2WikiMHQA while achieving competitive Faithfulness scores (78.35%–95.20%).
@@
CRAFT retains only <reason>+<answer>. CRAFT achieves the highest Faithfulness scores among all templates: 86.70% on MuSiQue, 96.00% on HotpotQA, and 94.10% on 2WikiMHQA. EM scores (56.91%, 66.51%, 79.39%) remain competitive, demonstrating that simplified templates can achieve strong performance with appropriate reward design.
@@
CRAFT uses direct answer generation without structured reasoning traces ( and not applicable). CRAFT achieves 56.25% EM on MuSiQue, 63.25% on HotpotQA, and 67.15% on 2WikiMHQA. A critical observation: CRAFT completely fails on this template (near-zero performance), indicating that smaller models require structured reasoning traces for stable training.
B.4 Format Compliance Analysis